# The algebra $\mathcal{D}(W)$ via strong Darboux transformations

### v1

@@ Abstract.
The Matrix Bochner Problem aims to classify weight matrices $W$ such that its algebra $\mathcal{D}(W)$, of all differential operators that have a sequence of these matrix orthogonal polynomials as eigenfunctions, contains a second-order differential operator. In [5] it is proven that, under certain assumptions, the solutions to the Matrix Bochner Problem can be obtained through a noncommutative bispectral Darboux transformation of some classical scalar weights.
@@ 1. Introduction
$\langle P\cdot\mathcal{V},Q\rangle_{\widetilde{W}}=\langle P,Q\cdot\mathcal{N}\rangle_{W},\quad\text{ for all }P,Q\in\operatorname{Mat}_{N}(\mathbb{C})[x].$
@@
To illustrate this deep connection between the algebras $\mathcal{D}(W)$ and $\mathcal{D}(\widetilde{W})$ when $W$ and $\widetilde{W}$ are strong Darboux transformations to each other, we consider several examples of irreducible weights $W$. We prove that they are a strong Darboux transformation of a direct sum of scalar weights, and thus we determine the algebra $\mathcal{D}(W)$ in a very simple way.
@@ 2. Matrix valued orthogonal polynomials and the algebra $\mathcal{D}(W)$
Let $W=W(x)$ be a weight matrix of size $N$ on the real line, that is, a complex $N\times N$ matrix-valued integrable function on the interval $(x_{0},x_{1})$ such that $W(x)$ is positive definite almost everywhere and with finite moments of all orders. Let $\operatorname{Mat}_{N}(\mathbb{C})$ be the algebra of all $N\times N$ complex matrices and let $\operatorname{Mat}_{N}(\mathbb{C})[x]$ be the algebra of polynomials in the indeterminate $x$ with coefficients in $\operatorname{Mat}_{N}(\mathbb{C})$. We consider the following Hermitian sesquilinear form in the linear space $\operatorname{Mat}_{N}(\mathbb{C})[x]$
@@
Given a weight matrix $W$ one can construct sequences $\{Q_{n}\}_{n\in\mathbb{N}_{0}}$ of matrix-valued orthogonal polynomials, i.e. the $Q_{n}$ are polynomials of degree $n$ with nonsingular leading coefficient and $\langle Q_{n},Q_{m}\rangle=0$ for $n\neq m$. We observe that there exists a unique sequence of monic orthogonal polynomials $\{P_{n}\}_{n\in\mathbb{N}_{0}}$ in $\operatorname{Mat}_{N}(\mathbb{C})[x]$. By following a standard argument (see [20] or [21]) one shows that the monic orthogonal polynomials $\{P_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy a three-term recursion relation
@@
Let $W=W(x)$ be a weight matrix of size $N$ and let $\{P_{n}\}_{n\geq 0}$ be the sequence of monic orthogonal polynomials in $\operatorname{Mat}_{N}(\mathbb{C})[x]$. If $D$ is differential operator of order $s$, as in (2), such that
@@
The formal adjoint on $\operatorname{Mat}_{N}(\Omega([[x]])$, denoted by $\mbox{}^{*}$, is the unique involution extending Hermitian conjugate on $\operatorname{Mat}_{N}(\mathbb{C}([x])$ and sending $\partial I$ to $-\partial I$. The formal $W$-adjoint of $\mathfrak{D}\in\operatorname{Mat}_{N}(\Omega([x])$, or the formal adjoint of $\mathfrak{D}$ with respect to $W(x)$ is the differential operator $\mathfrak{D}^{\dagger}\in\operatorname{Mat}_{N}(\Omega[[x]])$ defined by
@@
for all $P,Q\in\operatorname{Mat}_{N}(\mathbb{C})[x]$. Then we say that the operator $\widetilde{\mathfrak{D}}$ is the $W$-adjoint of $\mathfrak{D}$.
@@
To avoid pathological cases, we will consider only weight matrices which decrease exponentially at infinity. That is, weight matrices $W(x)$ such that there exist some positive real numbers $a,b$ such that
$\lim_{x\to\pm\infty}e^{a|x|^{b}}W(x)=0.$
With this in mind, the following proposition holds.
Proposition 2.4 ([5]).
If $\mathfrak{D}\in\operatorname{Mat}_{N}(\Omega[x])$ is $W$-adjointable and $\mathfrak{D}^{\dagger}\in\operatorname{Mat}_{N}(\Omega[x])$, then $\mathfrak{D}^{\dagger}$ is the $W$-adjoint of $\mathfrak{D}$, i.e.
$\langle\,P\mathfrak{D},Q\,\rangle=\langle\,P,\,Q{\mathfrak{D}}^{\dagger}\rangle,$
for all $P,Q\in\operatorname{Mat}_{N}(\mathbb{C})[x]$.
For $\mathfrak{D}=\sum_{j=0}^{n}\partial^{j}F_{j}\in\operatorname{Mat}_{N}(\Omega[x])$, the formal $W$-adjoint of $\mathfrak{D}$ is given by $\mathfrak{D}^{\dagger}=\sum_{k=0}^{n}\partial^{k}G_{k}$, with
@@
We say that a differential operator $D\in\mathcal{D}(W)$ is $W$-symmetric if $\langle P{D},Q\rangle=\langle P,QD\rangle$, for all $P,Q\in\operatorname{Mat}_{N}(\mathbb{C})[x]$. An operator $\mathfrak{D}\in\operatorname{Mat}_{N}(\Omega([x])$ is called formally $W$-symmetric if $\mathfrak{D}^{\dagger}=\mathfrak{D}$. In particular if $\mathfrak{D}\in\mathcal{D}(W)$, then $\mathfrak{D}$ is $W$-symmetric if and only if it is formally $W$-symmetric.
It is shown that the set $\mathcal{S}(W)$ of all symmetric operators in $\mathcal{D}(W)$ is a real form of the space $\mathcal{D}(W)$, i.e.
$\mathcal{D}(W)=\mathcal{S}(W)\oplus i\mathcal{S}(W),$
as real vector spaces.
@@
In particular, the coefficients of a differential operator of order two in $\mathcal{S}(W)$ satisfy the classical symmetry equations.
@@
We say that a differential operator $\mathcal{V}\in\operatorname{Mat}_{N}(\Omega)[x]$ is a degree-preserving operator if the degree of $P(x)\cdot\mathcal{V}$ is equal to the degree of $P(x)$ for all $P(x)\in\operatorname{Mat}_{N}(\mathbb{C})[x]$.
In particular, a degree-preserving differential operator $\mathcal{V}=\sum_{j=0}^{m}\partial^{j}F_{j}\in\operatorname{Mat}_{N}(\Omega)[x]$ satisfies $\deg(F_{j})\leq j$ for all $j$. If $P(x)\in\operatorname{Mat}_{N}(\mathbb{C})[x]$ has a nonsingular leading coefficient, then $P(x)\cdot\mathcal{V}$ also has a nonsingular leading coefficient. Furthermore, it should be noted that every degree-preserving differential operator is not a zero divisor in the algebra of differential operators $\operatorname{Mat}_{N}(\Omega)[x]$.
@@ 3. Strong Darboux transformation
Let $W$ and $\widetilde{W}$ be weight matrices supported on the same interval, and let $P_{n}(x)$ and $\widetilde{P}_{n}(x)$ be their associated sequence of monic orthogonal polynomials. We say that $\widetilde{W}$ is a Darboux transformation of $W$ if there exists a differential operator $D\in\mathcal{D}(W)$ that can be factorized as $D=\mathcal{V}\mathcal{N}$ with degree-preserving operators $\mathcal{V},\,\mathcal{N}\in\operatorname{Mat}_{N}(\Omega)[x]$, such that
@@
$\langle P\cdot\mathcal{V},Q\rangle_{\widetilde{W}}=\langle P,Q\cdot\mathcal{N}\rangle_{W}\qquad\text{ for all }P,Q\in\operatorname{Mat}_{N}(\mathbb{C})[x],$
@@
Let $P_{n}(x)$ and $\widetilde{P}_{n}(x)$ be the associated sequence of monic orthogonal polynomials for $W$ and $\widetilde{W}$ respectively. We need to show that $P_{n}(x)\cdot\mathcal{V}=A_{n}\widetilde{P}_{n}(x)$ for some sequence of matrices $A_{n}\in\operatorname{Mat}_{N}(\mathbb{C})$, and that $\mathcal{N}$ is a degree-preserving operator. Since $\mathcal{V}$ is degree-preserving, we have that $P_{n}(x)\cdot\mathcal{V}$ is a polynomial of degree $n$ with leading coefficient a nonsingular matrix. If $n\not=m$, then
@@
Therefore $P_{n}(x)\cdot\mathcal{V}$ is a sequence of orthogonal polynomials for $\widetilde{W}$. Thus, there exists a sequence of nonsingular matrices $A_{n}$ such that $A_{n}\widetilde{P}_{n}(x)=P_{n}(x)\cdot\mathcal{V}$. Finally, to demonstrate that $\mathcal{N}$ is a degree-preserving operator, consider $Q\in\operatorname{Mat}_{N}(\mathbb{C})[x]$ of degree $m$. We can express the polynomial $Q$ as $Q(x)=\sum_{j=0}^{m}K_{j}\widetilde{P}_{j}(x)$, where $K_{j}\in\operatorname{Mat}_{N}(\mathbb{C})$ and $K_{m}\neq 0$. Now, if $Q(x)\cdot\mathcal{N}$ is of degree lower than $m$, then
@@
$\langle P\cdot\mathcal{V},Q\rangle_{\widetilde{W}}=\langle P,Q\cdot\mathcal{N}\rangle_{W},\qquad\text{ for all }P,Q\in\operatorname{Mat}_{N}(\mathbb{C})[x].$
@@
Supose that $\widetilde{W}$ is a strong Darboux transformation of $W$, with strong Darboux transformer $\mathcal{V}$ and $\mathcal{N}=\widetilde{W}\mathcal{V}^{\ast}W^{-1}$ as in the definition. The leading coefficient of $\mathcal{N}$ is nonsingular because the one of $\mathcal{V}$ is. By Proposition 3.2 it follows that $\widetilde{D}=\mathcal{N}\mathcal{V}\in\mathcal{D}(\widetilde{W})$. Besides, $W\mathcal{N}^{\ast}\widetilde{W}^{-1}=\mathcal{V}$ and $\langle P\cdot\mathcal{N},Q\rangle_{W}=(\langle Q,P\cdot\mathcal{N}\rangle_{W})^{\ast}=(\langle Q\cdot\mathcal{V},P\rangle_{\widetilde{W}})^{\ast}=\langle P,Q\cdot\mathcal{V}\rangle_{\widetilde{W}}$ for all $P,Q\in\operatorname{Mat}_{N}(\mathbb{C})[x]$. Therefore $W$ is a strong Darboux transformation of $\widetilde{W}$ with strong Darboux transformer $\mathcal{N}$.
Finally, let $W_{1}$, $W_{2}$ and $W_{3}$ weights matrices such that $W_{1}\sim W_{2}$ and $W_{2}\sim W_{3}$. Let $\mathcal{V}_{1}$ be a strong Darboux transformer from $W_{1}$ to $W_{2}$ and $\mathcal{V}_{2}$ a strong Darboux transformer from $W_{2}$ to $W_{3}$. It is easy to verify that the differential operator $\mathcal{V}_{1}\mathcal{V}_{2}$ is a strong Darboux transformer from $W_{1}$ to $W_{3}$ and therefore $W_{1}\sim W_{3}$ and this completes the proof. ∎
@@
$W$ is a Darboux transformation of $w(x)=e^{-x^{2}}I$ but not a strong Darboux transformation of it.
@@
We have proved in [3] that the algebra $\mathcal{D}(w)$, for $w(x)=e^{-x^{2}}I$, is
$\mathcal{D}(w)=\sum_{j=1}^{3}\sum_{i=1}^{3}\mathbb{C}[\delta]E_{i,j},$
where $\delta=\partial^{2}+\partial(-2x)$ is the second-order differential operator of the classical scalar Hermite weight. Then, the operator
@@
The differential operators $\mathcal{V}$ and $\mathcal{N}$ are degree-preserving, and it follows that $\mathcal{N}=W\mathcal{V}^{\ast}w^{-1}$. The equality $\langle P\cdot\mathcal{V},Q\rangle_{W}=\langle P,Q\cdot\mathcal{N}\rangle_{w}$ holds for all $P,\,Q\in\operatorname{Mat}_{3}(\mathbb{C})[x]$ by integrating by parts since the boundary conditions will vanish due to the exponential decay of $e^{-x^{2}}$. Thus, by Proposition 3.4 we have that $W$ is a Darboux transformation of $e^{-x^{2}}I$.
Now, we assume that $W$ is a strong Darboux transformation $w(x)=e^{-x^{2}}I$. Since $\mathcal{D}(e^{-x^{2}}I)$ contains only even-order differential operators, we have, by Corollary 3.11, that $\mathcal{D}(W)$ must contain only even-order differential operators. However, by Proposition 3.3, we have that $\mathcal{A}=\mathcal{N}\begin{pmatrix}0&&0&&0\\ 0&&1&&0\\ 0&&0&&0\end{pmatrix}\mathcal{V}\in\mathcal{D}(W)$ and is of odd order, which is a contradiction. ∎
@@
$Q_{n}(x)=H_{n}(x)\cdot\mathcal{V}=\begin{pmatrix}-\frac{2}{a}H_{n}(x)&&H_{n}^{\prime}(x)&&-bxH_{n}^{\prime}(x)\\ H_{n}^{\prime\prime}(x)-\frac{4}{ab}H_{n}(x)&&-axH_{n}^{\prime\prime}(x)&&abx^{2}H_{n}^{\prime\prime}(x)+\frac{4}{ab}H_{n}(x)\\ -\frac{4}{ab}H_{n}^{\prime}(x)&&\frac{4x}{b}H_{n}^{\prime}(x)+\frac{8}{ba^{2}}H_{n}(x)&&-4xH_{n}^{\prime\prime}(x)-\frac{2(a^{2}+2)}{a^{2}}H_{n}^{\prime}(x)\end{pmatrix},$

### v2

@@ Abstract.
The Matrix Bochner Problem aims to classify weight matrices $W$ such that the algebra $\mathcal{D}(W)$, of all differential operators that have a sequence of matrix-valued orthogonal polynomials for $W$ as eigenfunctions, contains a second-order differential operator. In [5] it is proven that, under certain assumptions, the solutions to the Matrix Bochner Problem can be obtained through a noncommutative bispectral Darboux transformation of some classical scalar weights.
@@ 1. Introduction
$\langle P\cdot\mathcal{V},Q\rangle_{\widetilde{W}}=\langle P,Q\cdot\mathcal{N}\rangle_{W},\quad\text{ for all }P,Q\in\operatorname{Mat}_{N}(\mathbb{C}[x]).$
@@
To illustrate this deep connection between the algebras $\mathcal{D}(W)$ and $\mathcal{D}(\widetilde{W})$ when $W$ and $\widetilde{W}$ are strong Darboux transformations of each other, we consider several examples of irreducible weights $W$. We prove that they are a strong Darboux transformation of a direct sum of scalar weights, and thus we determine the algebra $\mathcal{D}(W)$ in a very simple way.
@@ 2. Matrix valued orthogonal polynomials and the algebra $\mathcal{D}(W)$
Let $W=W(x)$ be a weight matrix of size $N$ on the real line, that is, a complex $N\times N$ matrix-valued smooth function on the interval $(x_{0},x_{1})$ such that $W(x)$ is positive definite almost everywhere and with finite moments of all orders. Let $\operatorname{Mat}_{N}(\mathbb{C})$ be the algebra of all $N\times N$ complex matrices and let $\operatorname{Mat}_{N}(\mathbb{C}[x])$ be the algebra of polynomials in the indeterminate $x$ with coefficients in $\operatorname{Mat}_{N}(\mathbb{C})$. We consider the following Hermitian sesquilinear form in the linear space $\operatorname{Mat}_{N}(\mathbb{C}[x])$
@@
Given a weight matrix $W$ one can construct sequences $\{Q_{n}\}_{n\in\mathbb{N}_{0}}$ of matrix-valued orthogonal polynomials, i.e. the $Q_{n}$ are polynomials of degree $n$ with nonsingular leading coefficient and $\langle Q_{n},Q_{m}\rangle=0$ for $n\neq m$. We observe that there exists a unique sequence of monic orthogonal polynomials $\{P_{n}\}_{n\in\mathbb{N}_{0}}$ in $\operatorname{Mat}_{N}(\mathbb{C}[x])$. By following a standard argument (see [20] or [21]) one shows that the monic orthogonal polynomials $\{P_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy a three-term recursion relation
@@
Let $W=W(x)$ be a weight matrix of size $N$ and let $\{P_{n}\}_{n\geq 0}$ be the sequence of monic orthogonal polynomials in $\operatorname{Mat}_{N}(\mathbb{C}[x])$. If $D$ is differential operator of order $s$, as in (2), such that
@@
The formal adjoint on $\operatorname{Mat}_{N}(\Omega[[x]])$, denoted by $\mbox{}^{*}$, is the unique involution extending Hermitian conjugate on $\operatorname{Mat}_{N}(\mathbb{C}[x])$ and sending $\partial I$ to $-\partial I$. The formal $W$-adjoint of $\mathfrak{D}\in\operatorname{Mat}_{N}(\Omega[x])$, or the formal adjoint of $\mathfrak{D}$ with respect to $W(x)$ is the differential operator $\mathfrak{D}^{\dagger}\in\operatorname{Mat}_{N}(\Omega[[x]])$ defined by
@@
for all $P,Q\in\operatorname{Mat}_{N}(\mathbb{C}[x])$. Then we say that the operator $\widetilde{\mathfrak{D}}$ is the $W$-adjoint of $\mathfrak{D}$.
@@
In this work, to avoid pathological weights and their technicalities, we consider only weight matrices that are “good enough”, i.e., such that for each integer $n\geq 0$ the $n$-th derivative $W^{(n)}(x)$ decreases exponentially at infinity and there exists a scalar polynomial $p_{n}(x)$ such that $W^{(n)}(x)p_{n}(x)$ has finite moments. See [5, Section 2.2].
With the above remark in mind, it follows that the $W$-adjoint of any $W$-adjointable operator in $\operatorname{Mat}_{N}(\Omega[x])$ coincides with the formal $W$-adjoint. For $\mathfrak{D}=\sum_{j=0}^{n}\partial^{j}F_{j}\in\operatorname{Mat}_{N}(\Omega[x])$, the formal $W$-adjoint of $\mathfrak{D}$ is given by $\mathfrak{D}^{\dagger}=\sum_{k=0}^{n}\partial^{k}G_{k}$, with
@@
We say that a differential operator $D\in\mathcal{D}(W)$ is $W$-symmetric if $\langle P{D},Q\rangle=\langle P,QD\rangle$, for all $P,Q\in\operatorname{Mat}_{N}(\mathbb{C}[x])$. In particular, since all operators in $\mathcal{D}(W)$ are $W$-adjointable, we have that $D\in\mathcal{D}(W)$ is $W$-symmetric if and only if $D=D^{\dagger}$.
@@
In particular, the coefficients of a differential operator of order two in $\mathcal{S}(W)$, the set of all $W$-symmetric operators in $\mathcal{D}(W)$, satisfy the classical symmetry equations:
@@
On the other hand, in [18, Corollary 4.5] it is proved that the algebra $\mathcal{D}(W)$ satisfies
$\mathcal{D}(W)=\mathcal{S}(W)\oplus i\mathcal{S}(W).$
@@
We say that a differential operator $\mathcal{V}\in\operatorname{Mat}_{N}(\Omega[x])$ is a degree-preserving operator if the degree of $P(x)\cdot\mathcal{V}$ is equal to the degree of $P(x)$ for all $P(x)\in\operatorname{Mat}_{N}(\mathbb{C}[x])$.
In particular, a degree-preserving differential operator $\mathcal{V}=\sum_{j=0}^{m}\partial^{j}F_{j}\in\operatorname{Mat}_{N}(\Omega[x])$ satisfies $\deg(F_{j})\leq j$ for all $j$. If $P(x)\in\operatorname{Mat}_{N}(\mathbb{C}[x])$ has a nonsingular leading coefficient, then $P(x)\cdot\mathcal{V}$ also has a nonsingular leading coefficient. Furthermore, it should be noted that every degree-preserving differential operator is not a zero divisor in the algebra of differential operators $\operatorname{Mat}_{N}(\Omega[x])$.
@@ 3. Strong Darboux transformation
Let $W$ and $\widetilde{W}$ be weight matrices supported on the same interval, and let $P_{n}(x)$ and $\widetilde{P}_{n}(x)$ be their associated sequence of monic orthogonal polynomials. We say that $\widetilde{W}$ is a Darboux transformation of $W$ if there exists a differential operator $D\in\mathcal{D}(W)$ that can be factorized as $D=\mathcal{V}\mathcal{N}$ with degree-preserving operators $\mathcal{V},\,\mathcal{N}\in\operatorname{Mat}_{N}(\Omega[x])$, such that
@@
$\langle P\cdot\mathcal{V},Q\rangle_{\widetilde{W}}=\langle P,Q\cdot\mathcal{N}\rangle_{W}\qquad\text{ for all }P,Q\in\operatorname{Mat}_{N}(\mathbb{C}[x]),$
@@
Let $P_{n}(x)$ and $\widetilde{P}_{n}(x)$ be the associated sequence of monic orthogonal polynomials for $W$ and $\widetilde{W}$ respectively. We need to show that $P_{n}(x)\cdot\mathcal{V}=A_{n}\widetilde{P}_{n}(x)$ for some sequence of matrices $A_{n}\in\operatorname{Mat}_{N}(\mathbb{C})$, and that $\mathcal{N}$ is a degree-preserving operator. Since $\mathcal{V}$ is degree-preserving, we have that $P_{n}(x)\cdot\mathcal{V}$ is a polynomial of degree $n$ whose leading coefficient is a nonsingular matrix. If $n\not=m$, then
@@
Therefore $P_{n}(x)\cdot\mathcal{V}$ is a sequence of orthogonal polynomials for $\widetilde{W}$. Thus, there exists a sequence of nonsingular matrices $A_{n}$ such that $A_{n}\widetilde{P}_{n}(x)=P_{n}(x)\cdot\mathcal{V}$. Finally, to demonstrate that $\mathcal{N}$ is a degree-preserving operator, consider $Q\in\operatorname{Mat}_{N}(\mathbb{C}[x])$ of degree $m$. We can express the polynomial $Q$ as $Q(x)=\sum_{j=0}^{m}K_{j}\widetilde{P}_{j}(x)$, where $K_{j}\in\operatorname{Mat}_{N}(\mathbb{C})$ and $K_{m}\neq 0$. Now, if $Q(x)\cdot\mathcal{N}$ is of degree lower than $m$, then
@@
$\langle P\cdot\mathcal{V},Q\rangle_{\widetilde{W}}=\langle P,Q\cdot\mathcal{N}\rangle_{W},\qquad\text{ for all }P,Q\in\operatorname{Mat}_{N}(\mathbb{C}[x]).$
@@
Suppose that $\widetilde{W}$ is a strong Darboux transformation of $W$, with strong Darboux transformer $\mathcal{V}$ and $\mathcal{N}=\widetilde{W}\mathcal{V}^{\ast}W^{-1}$ as in the definition. The leading coefficient of $\mathcal{N}$ is nonsingular because the one of $\mathcal{V}$ is. By Proposition 3.2 it follows that $\widetilde{D}=\mathcal{N}\mathcal{V}\in\mathcal{D}(\widetilde{W})$. Besides, $W\mathcal{N}^{\ast}\widetilde{W}^{-1}=\mathcal{V}$ and $\langle P\cdot\mathcal{N},Q\rangle_{W}=(\langle Q,P\cdot\mathcal{N}\rangle_{W})^{\ast}=(\langle Q\cdot\mathcal{V},P\rangle_{\widetilde{W}})^{\ast}=\langle P,Q\cdot\mathcal{V}\rangle_{\widetilde{W}}$ for all $P,Q\in\operatorname{Mat}_{N}(\mathbb{C}[x])$. Therefore $W$ is a strong Darboux transformation of $\widetilde{W}$ with strong Darboux transformer $\mathcal{N}$.
Finally, let $W_{1}$, $W_{2}$ and $W_{3}$ be weight matrices such that $W_{1}\sim W_{2}$ and $W_{2}\sim W_{3}$. Let $\mathcal{V}_{1}$ be a strong Darboux transformer from $W_{1}$ to $W_{2}$ and $\mathcal{V}_{2}$ a strong Darboux transformer from $W_{2}$ to $W_{3}$. It is easy to verify that the differential operator $\mathcal{V}_{1}\mathcal{V}_{2}$ is a strong Darboux transformer from $W_{1}$ to $W_{3}$ and therefore $W_{1}\sim W_{3}$ and this completes the proof. ∎
@@
$W$ is a Darboux transformation of $\widetilde{W}(x)=e^{-x^{2}}I$ but not a strong Darboux transformation of it.
@@
We have proved in [3] that the algebra $\mathcal{D}(\widetilde{W})$, for $\widetilde{W}(x)=e^{-x^{2}}I$, is
$\mathcal{D}(\widetilde{W})=\sum_{j=1}^{3}\sum_{i=1}^{3}\mathbb{C}[\delta]E_{i,j},$
where $\delta=\partial^{2}+\partial(-2x)$ is the second-order differential operator of the classical scalar Hermite weight, and $E_{i,j}$ denotes the matrix with zeros everywhere except for a $1$ in the $(i,j)$-entry. Then, the operator
@@
The differential operators $\mathcal{V}$ and $\mathcal{N}$ are degree-preserving, and it follows that $\mathcal{N}=W\mathcal{V}^{\ast}\widetilde{W}^{-1}$. The equality $\langle P\cdot\mathcal{V},Q\rangle_{W}=\langle P,Q\cdot\mathcal{N}\rangle_{\widetilde{W}}$ holds for all $P,\,Q\in\operatorname{Mat}_{3}(\mathbb{C}[x])$ by integrating by parts since the boundary conditions will vanish due to the exponential decay of $e^{-x^{2}}$. Thus, by Proposition 3.4 we have that $W$ is a Darboux transformation of $e^{-x^{2}}I$.
Now, we assume that $W$ is a strong Darboux transformation $\widetilde{W}(x)=e^{-x^{2}}I$. Since $\mathcal{D}(e^{-x^{2}}I)$ contains only even-order differential operators, we have, by Corollary 3.11, that $\mathcal{D}(W)$ must contain only even-order differential operators. However, by Proposition 3.3, we have that $\mathcal{A}=\mathcal{N}\begin{pmatrix}0&&0&&0\\ 0&&1&&0\\ 0&&0&&0\end{pmatrix}\mathcal{V}\in\mathcal{D}(W)$ and is of odd order, which is a contradiction. ∎
@@
$Q_{n}(x)=H_{n}(x)\cdot\mathcal{V}=\begin{pmatrix}-\frac{2}{a}H_{n}(x)&H_{n}^{\prime}(x)&-bxH_{n}^{\prime}(x)\\ H_{n}^{\prime\prime}(x)-\frac{4}{ab}H_{n}(x)&-axH_{n}^{\prime\prime}(x)&abx^{2}H_{n}^{\prime\prime}(x)+\frac{4}{ab}H_{n}(x)\\ -\frac{4}{ab}H_{n}^{\prime}(x)&\frac{4x}{b}H_{n}^{\prime}(x)+\frac{8}{ba^{2}}H_{n}(x)&H_{n}(x)^{\prime\prime\prime}-4xH_{n}^{\prime\prime}(x)-\frac{2(a^{2}+2)}{a^{2}}H_{n}^{\prime}(x)\end{pmatrix},$